Hi, I am in need of help! if you can help me solve that would be great!thank you
The mean gas mileage for a hybrid car is 57 miles per gallon. Suppose that the gasoline mileage is approximately normally distributed with a standard deviation of 3.5 miles per gallon.​
a) What proportion of hybrids gets over 62 miles per​ gallon?
b) What proportion of hybrids gets 53 miles per gallon or​ less?
c) what proportion of hybrids gets between 59 and 62 miles per​ gallon?
d) What is the probability that a randomly selected hybrid gets less than 46 miles per​ gallon?

Answer by Theo(13342)

you can use a calculator or you can use the normal distribution tables.

first step is to get the z-scores.

the formula for z-score is:

z = (x - m) / s

z is the z-score
x i the raw score
m is the mean
s is the standard deviation

you are given that the mean is 57 and the standard deviation is 3.5

if the distribution is normal, then the mean, median, and mode coincide at the middle of the distribution.

the mean is the average
the median is the point where half of the data falls below and half of the data falls above.
the mode is is the most frequently occurring value.

here's a description.
https://statistics.laerd.com/statistical-guides/measures-central-tendency-mean-mode-median.php

the total area underneath the normal distribution curve is equal to 1.
this id done by design.
a point, on the normal distribution curve has a certain amount of the area to the left of it and to the right of it.

this area is expressed as a proportion or percent.

for example, the mean is exactly in the middle and, since the curve is symmetric, half the area is to the left and half the area is to the right.
the right.

the z-score at the mean = 0.
all z-scores are in relation to the mean.
z-scores to the left of the mean are negative.
z-scores to the right of the mean are positive.

the tables are set up to give you the area to the left of the z-score.
if you want the area to the right of the z-score, you would subtract the area from 1.

for example:
a z-score of 1 gives you an area to the left of it as .84134 to the left of it.
the area to the right of it would be equal to 1 minus .84134 = .15866.

since the z-scores are symmetric about the mean, then a z-score of - 1 would have an area of .15866 to the left of it and an area of 1 - .15866 = .84134 to the right of it.

the sum of the areas to the left of the z-score and to the right of the z-scores will always be equal to 1.

to use the z-score tables, you need to translate the raw score to the z-score.




here's a table where you can check that out for yourself.

https://www.math.arizona.edu/~rsims/ma464/standardnormaltable.pdf

the z-score formula was mentioned above and is repeated here.

the formula is z = (x - m) / s

x is the raw score
m is the mean
s is the standard deviation.

you are given that the mean is 57 and the standard deviation is 3.5

your questions are below and will be answedred in turn.

a) What proportion of hybrids gets over 62 miles per​ gallon?

z = (62 - 57) / 3.5
solve for z to get:
z = 5/3.5 = 1.43 rounded to 2 decimal places.
look in the table to get area to the left of that z-score = .92364.
area to the right of it = 1 - .92364 = .07636.
you are looking for the area to the right of it, therefore:
the proportion of hybrids that get over 62 miles per gallon is .07636.

b) What proportion of hybrids gets 53 miles per gallon or​ less?

z = (53 - 57) / 3.5
solve for z to get:
z = -4/3.5 = -1.14 rounded to 2 decimal places.
look in the table to get area to the left of that z-score = .12714.
area to the right of it = 1 - .12714 = .87286.
you are looking for the area to the left of it, therefore:
the proportion of hybrids that get less than 53 miles per gallon is .12714.

c) what proportion of hybrids gets between 59 and 62 miles per​ gallon?

you need the low z-score and the high z-score.
for the low z-score:
z = (59 - 57) / 3.5
solve for z to get:
z = 0.57 rounded to 2 decimal places.
look in the table to get area to the left of that = .71566.
area to the right of it = 1 - .71566 = .28434

for the high z-score:
z = (62 - 57) / 3.5
solve for z to get:
z = 5/3.5 = 1.43 rounded to 2 decimal places.
look in the table to get area to the left of that = .92364.
area to the right of that = 1 - .92364 = .07636.

for this problem, you need the area to the left of both z-scores.
you would subtract the smaller area from the larger area to get:
.92364 - .71566 = .20798.
the proportion of hybrids that get between 53 and 62 miles per gallon is .20798.

d) What is the probability that a randomly selected hybrid gets less than 46 miles per​ gallon?

z = (46 - 57) / 3.5
solve for z to get:
z = -11/3.5 = -3.14 rounded to 2 decimal places.
look in the table to get area to the left of that = .00084
area to the right of that = 1 - .00084 = .99916
you want the area to the left of that.

the proportion of cars that get less than 46 miles per gallon is .00084.

that is also the probability that a randomly selected car will get less than 46 miles per gallon.

that should answer all your questions.
some additional comments below.

proportion * 100 = percent
so if the proportion of cars that get less than 46 miles per gallon is .00084, then the percent of cars that get less than 46 miles per gallon is .084%.

proportions and percents and probability are related.
if the proportion of cars that gets less than 62 miles per gallon is .92364, then the percent of cars that get less than 62 miles per gallon is 92.364% and the probability that a randomly selected car will get less than 62 miles per gallon is .92364.

the tables are set up to give you the area to the left of the z-score.
you look in the first column to find the z-score to the first decimal digit.
you then look across the rows to get the z-score to the second decimal digit.

for example, to find the area to the left of a z-score of -3.56, you look down the first column to find a z-score of -3.5.
you then look across that row to find the z-score under .06.
the area that you find at the intersection of the row that -3.5 is on and the column that .06 is on is the area to the left of the z-score of -3.56.

here's a picture of what i mean.



the area to the left of a z-score of -3.56 is equal to .00019.

so far we've use the tables.
however, there are calculators on the web that can help with this.
one such calculator can be found at https://www.calculator.net/z-score-calculator.html?c1raw=3.075&c1mean=3.2&c1sd=.1&calctype=zscore&x=81&y=18

for example, area to the left of a z-score of 1.43 would be shown as .92364 in this calculator.
this agrees with the number you got from the table.
if you entered the raw score rather than the z-score, the calculator would have told you that the area to the left of the raw score was equal to .92344 and the z-score was 1.42857.

there are 3 sections to the table.
learn to use those sections and they can help you as well.
the page the calculator is on also shows you the z-score table.

all in all, a pretty decent piece of work that can help you once you learn how to use it properly.

here's a couple of results from the examples i used.






if you have any questions about any of this, just let mem know.

theo